Solve the following PDE using the method of separation of variables and Fourier series

utt = c2uxx, 0<x<1, t<0

u(0,t) = 0, t>0

u(1,t) = 0, t>0

u(x,0) = 2x(1-x), 0<=x<=1

ut(x,0) =0


1
Expert's answer
2021-06-28T03:56:09-0400

We consider the wave equation satisfying Dirichlet boundary condition:

"u_{tt}=c^2u_{xx}" , "0<x<1,\ \ t>0"

"u(0,t)=u(1,t)=0" , "t\geq 0"

"u(x,0)=2x(1-x),\quad 0\leq x\leq 1"

"u_t(x,0)=0,\quad\quad \quad \quad \ 0\leq x\leq 1"


We look for solutions of the form "u(x,t)=\sum \limits_{n=1}^\infty\big(A_n\cos (\pi nct)+B_n\sin (\pi nct)\big) \sin (\pi nx)"

Setting "t=0" , we get

"u(x,0)= \sum \limits_{n=1}^\infty A_n\sin (\pi nx)=2x(1-x)"

Coefficients "A_n" are equal to "A_n=2\int \limits_0^1 2x(1-x)\sin (\pi nx)dx=- \frac{\cos (\pi nx)}{\pi n} 4x(1-x)\bigg|_0^1+\int\limits_0^1 \frac{\cos (\pi nx)}{\pi n}4(1-2x)dx= \int\limits_0^1 \frac{\cos (\pi nx)}{\pi n}4(1-2x)dx= \frac{\sin (\pi nx)}{\pi ^2 n^2 }4(1-2x)\bigg|_0^1 +\frac{8}{\pi^2n^2}\int\limits_0^1 \sin (\pi n x)dx=-\frac{\cos (\pi n x)}{\pi n}\cdot \frac{8}{\pi^2n^2}\bigg|_0^1 =\frac{8}{\pi^3n^3}(1-\cos(\pi n))=\begin{cases}
0, \quad n=2k
\\
\frac{16}{\pi^3n^3}, \quad n=2k+1
\end{cases}"



"u_t(x,t)=\sum \limits_{n=1}^\infty \pi n c\big(B_n\cos (\pi nct)-A_n\sin (\pi nct)\big) \sin (\pi nx)"

Setting "t=0" , we get 

"u_t(x,0)=\sum\limits_{n=1}^\infty \pi n cB_n\sin (\pi n x)=0"

Therefore, "B_n=0\ \ \forall n" .


Answer: "u(x,t)=\sum\limits_{k=0}^\infty \frac{16}{\pi^3(2k+1)^3}\cos (\pi (2k+1)ct) \sin (\pi(2k+1)x)."